# Diffusion-Driven Universal Model Inversion Attack for Face Recognition

### v1

@@ II-D2 Face Detection via MTCNN
II-E Step (b) - Top N Latent Code Selection
In Step (a), DiffUMI generates a set of $V$ randomly sampled yet reliable latent codes and their corresponding reconstructions $\{(x_{G_{v}},\hat{x}_{v})\}_{v=1}^{V}$. Instead of processing all $V$ candidates, it may select the one with the highest initial embedding similarity to the target, expecting lowest computational costs and distortions. However, the latent code yielding the best initial reconstruction may not always produce the optimal result after adversarial manipulation (Fig. 6). To address this, DiffUMI refines selection by choosing the top $N$ latent codes from $V$ (Fig. 15 of Appendix -D), where increasing $N$ improves model inversion performance but incurs higher computational overhead.
@@ II-E Step (b) - Top N Latent Code Selection
In particular, the $V$ reliable pairs are input into the target face recognition model, which outputs $V$ feature embeddings. Note that $\hat{x}_{v}=G(x_{G_{v}})$, previously computed and stored during Step (a) for face detection (Sec. II-D2), is used. Similarities, as described in Eq. 6, are calculated between these $V$ embeddings and the target embedding. The top $N$ embeddings, exhibiting the highest similarity, are retained for Step (c) as the initial latent codes for adversarial manipulation.
In the black-box setting, as outlined in Sec. II-B, each of the $(x_{G_{v}},\hat{x}_{v})$ pairs requires access to the target model, leading to a total of $V$ queries in Step (b):
@@
II-F Step (c) - Latent Code Manipulation
@@
As outlined in Sec. II-C, a randomly sampled latent code rarely satisfies the attack objective of reconstructing a facial image that shares the same identity as the target, even for the top-1 selection from Step (b). Therefore, refining the initial latent code is crucial to align the reconstructed face with the target. Existing methods manipulate latent codes starting from either the top-1 code (as discussed in Sec. II-E) or a latent code fusion of the top $N$ selections [8]. However, we observe that fusing reduces similarity significantly compared to the top-1 or top-2 selections. Additionally, relying on a single fused initialization may fail to ensure effective reconstruction, akin to the limitations of the top-1 approach. To overcome these challenges, we propose the Ranked Adversary method, which retains the top $N$ selection strategy while prioritizing latent codes with higher similarity for earlier adversarial manipulation. This approach introduces early termination when the attack objective is met, optimizing computational efficiency while preserving effectiveness.
II-F1 Algorithm of Ranked Adversary
@@
The Ranked Adversary approach (Fig. 16 of Appendix -D) initiates the refinement process with the top-1 latent code $x_{G_{1}}$, applying an adversarial attack strategy to iteratively adjust $x_{G_{1}}$ in pursuit of the objective defined in Eq. 7, guided by the objective function in Eq. 6. The process concludes when the similarity measure $\mathcal{L}_{1}$ exceeds the predefined attack threshold $\tau_{A}$, indicating successful manipulation where the reconstructed image sufficiently matches the target. In such cases, the reconstructed image $\hat{x}_{1}=G(x_{G_{1}}+\delta_{1})$ is produced as the final model inversion result. If $\mathcal{L}_{1}$ remains below $\tau_{A}$ after the maximum adversarial iterations $t_{max}$, the top-2 latent code $x_{G_{2}}$ undergoes the same optimization process. If none of the top $N$ latent codes achieve $\mathcal{L}_{n}\geq\tau_{A},\ n=1,\dots,N$, the reconstruction with the highest $\mathcal{L}_{n}$ is selected as the final output.
@@
As shown in Fig. 7, setting $\tau_{A}=\tau_{F}$ may ensure successful target matching, but only on the target model. Increasing $\tau_{A}$ beyond $\tau_{F}$ enhances robustness, particularly on the test model. In contrast, excessively high $\tau_{A}$ or the absence of early stopping may lead to the attack proceeds beyond the point of achieving $\mathcal{L}>0.98$, yielding negligible gains in optimization while exacerbating overfitting and diminishing generalization to the test model, as achieving such a threshold often forces adversarial manipulation to introduce artifacts that overfit the target model and exhaust all attack iterations. Therefore, we define $\tau_{A}$ for attacking the specific model as the maximum embedding similarity achievable by real facial images in this model, which represents the best case where, ideally, without any overfitting artifacts, the model can achieve the best similarity within the same identity. To compute this, we feed the entire face dataset $\mathcal{X}_{real}$ (all real faces) into the target model, then compute the maximum similarity between any two images. The similarity between different identities is usually lower than that between the same identity. Thus, we define $\tau_{A}$ as follows:
@@
II-F2 Adversarial Attack
Ranked Adversary employs fine-grained adversarial attack algorithms, as defined by Wang et al. [33], utilizing APGD [32] in the white-box setting and GreedyPixel [33] in the black-box setting. While various adversarial attack algorithms can be adapted for latent code manipulation within our framework, our empirical observation shows that only these fine-grained methods yield satisfactory reconstructions without introducing significant distortions or artifacts. This is due to the sensitivity of the diffusion model’s latent space, where even small perturbations can notably degrade reconstruction quality.
Additionally, we evaluated two alternative black-box attacks: Square attack [44], which uses the same $L_{2}$-norm constrained perturbation magnitude $\epsilon$ as APGD, and BruSLe attack [45], a pixel-wise attack like GreedyPixel that enforces sparsity constraints rather than directly constraining $\epsilon$. However, both methods performed inferiorly compared to GreedyPixel, as detailed in Sec. V-B.
II-F3 Query Efficiency in the Black-Box Setting
@@
III Out-Of-Domain Detection
Deep learning models map both in-domain and out-of-domain inputs to the same feature space, making it challenging to distinguish between them based solely on embeddings. For example, in a face recognition system, both a human face and a non-face object (e.g., a cat) produce embeddings with identical dimensions and numerical ranges, despite their semantic differences. To address this, we leverage the high-fidelity reconstruction capabilities of our model inversion framework to develop an application of OODD that identifies out-of-domain inputs based on their embeddings.
As illustrated in Fig. 8, two common failure cases arise in model inversion, sometimes concurrently. In-domain inputs, such as facial images, typically yield accurate, recognizable reconstructions. In contrast, out-of-domain inputs are more likely to exhibit at least one of these failure cases. Hence, we classify an input as out-of-domain if either failure is detected.
@@ III Out-Of-Domain Detection
where $\tau_{F_{m}}$ is the predefined similarity threshold, corresponding to the minimum equal error rate in face recognition.
@@
  • 1 A higher detection confidence threshold is applied during latent code generation to ensure reliable initialization, as generating a random Gaussian template is computationally efficient.
  • Refer to Tab. XIII in Appendix -C for notation definitions.
@@
IV Experimental Settings
Face Recognition Models. We assess model inversion attacks on two widely used face recognition models, FaceNet [36] and ArcFace [37], and two privacy-preserving models, DCTDP [5] and PartialFace [6], which are designed to mitigate inversion attacks. The decision thresholds for all models are set at the minimum equal error rate, as outlined in Tab. II.
@@ IV Experimental Settings
Benchmark. We benchmark DiffUMI against MAP2V [8] in the white-box setting. As shown in Tab. I, MAP2V is the most recent and most relevant state-of-the-art baseline, sharing our training-free setup and focus on open-set face recognition. While DiffUMI also outperforms training-dependent attacks [22, 12], as detailed in Appendix -A, they are not our primary baseline due to their limitation that each target model requires a separately trained generator, resulting in substantial computational overhead. We exclude closed-set face recognition attacks, as they rely on class labels rather than target embeddings, reflecting fundamentally different assumptions.
Experimental Setup. The configuration details for the white-box DiffUMI attack and OODD are provided in Tab. II, while the settings for the black-box DiffUMI attack are outlined in Fig. 10 of Sec. V-B. Specifically, $\tau_{F}$ is the cosine similarity threshold in face recognition, optimized to minimize the equal error rate. $\tau_{K}$ and $\tau_{D}$ in latent code generation are set close to the upper bound (1.0) to ensure highly reliable initialization. $\tau_{A}$ is defined in Eq. 14. For OODD, $\tau_{D}$ is set to maintain a maximum false positive rate of 5% for in-domain inputs. In the ablation studies, we vary one parameter at a time while holding all others constant. All experiments are conducted on an NVIDIA A100 40GB GPU.
Evaluation Protocols. Model inversion attacks and OODD are executed directly on the target model, often leading to outputs that overfit this model. Consequently, exceptional performance on the target model may stem from adversarial artifacts rather than the accurate reconstruction of facial features, a scenario classified as a failure in Sec. III. To thoroughly evaluate the effectiveness of privacy attacks and OODD, we adopt the following evaluation protocols.
@@
[…]
$\quad\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}\mathbb{I}\left(S(F(\hat{x_{i}}),F(x^{j\neq T}_{i}))\geq\tau_{F}\right)}{I},$
@@
For further validation, we also report Type I accuracy results in Appendix -E, which reflect the strength of the inversion attack and highlight potential privacy vulnerabilities.
OODD Rate: The OODD rate quantifies the failure rate of model inversion, based on the two failure cases outlined in Sec. III. For in-domain inputs, a lower OODD rate is preferable, indicating fewer inversion failures. Conversely, for out-of-domain inputs, a higher OODD rate is desirable, reflecting improved detection performance.
Joint Evaluation Using Target and Test Models: Test models, as defined in Sec. III, refer to those that differ from the target model. Evaluating on these models helps mitigate overfitting to the target model. In this study, four face recognition models are utilized, with one designated as the target model and the remaining three as test models for each attack. However, testing exclusively on test models may introduce bias, as the test models vary for each target model, leading to an unfair comparison of model vulnerability to privacy risks.
To address this, we adopt a joint evaluation approach that includes both the target and test models, ensuring consistent evaluation across all models. This approach mitigates the impact of overfitting to the target model alone.
Note that OODD is evaluated exclusively on test models, as detailed in Sec. III. This is because out-of-domain inputs can only be matched with their reconstruction (similar to Type I accuracy), and such matching is always successful on the target model due to overfitting.
V Performance
V-A White-Box Model Inverison
We assess the vulnerability of face recognition models to privacy threats using DiffUMI in the white-box setting. As shown in Tab. III, all four models, including privacy-preserving variants, fail to prevent privacy leakage. DiffUMI achieves Type II accuracy between 74.51% and 94.72% across all models and datasets. Notably, the oldest standard model, FaceNet, demonstrates the highest resistance, highlighting the limitations of existing privacy-preserving techniques.
@@ V-A White-Box Model Inverison
Compared to the benchmark, as shown in Tab. IV, DiffUMI consistently surpasses MAP2V in Type II accuracy and achieves similarity values closer to the attack threshold $\tau_{A}$ across all scenarios. The “random” rows represent unguided diffusion model generations, which fail to match targets, emphasizing the necessity of a structured inversion approach. The visualization in Fig. 9 further highlights DiffUMI’s superiority, demonstrating higher identity recovery accuracy than MAP2V. Our approach produces full headshot-style reconstructions, achieving optimal granularity for identity recovery.
TABLE IV: Performance of our DiffUMI across four face recognition models, compared to the benchmark attack MAP2V [8].
@@
  • 3 Average Type II accuracy across four test models [37, 36, 5, 6], as per the joint evaluation using target and test models outlined in Sec. IV.
@@
Figure 10: Performance of DiffUMI in the Black-Box Setting. Black-box DiffUMI achieves comparable reconstruction accuracy to its white-box counterpart (columns 3 vs. 10) in terms of Type II accuracy and OODD rate but incurs significantly higher computational costs in query complexity and runtime. Among black-box attack methods (columns 4-6), the fine-grained GreedyPixel algorithm, which introduces adversarial patterns akin to white-box attacks, is the only viable approach for manipulating diffusion model latent codes. Comparing columns 6–10, increasing the attack budget (query or sparsity) enhances model inversion accuracy but significantly raises computational costs. The target model is the privacy-preserving PartialFace [6], and the test model is ArcFace [37].
V-B Black-Box Model Inversion
We evaluate DiffUMI in the black-box setting using various adversarial attack algorithms and compare its performance to the white-box counterpart. As shown in Fig. 10 (columns 3 vs. 10), black-box DiffUMI (GreedyPixel) achieves Type II accuracy comparable to its white-box version (APGD) but incurs substantially higher computational costs in time and queries. Due to these constraints, evaluations are conducted on only ten LFW samples [35] using the privacy-preserving PartialFace model [6], with efficiency improvements left for future work. Moreover, Fig. 10 (columns 4–6) demonstrates that only fine-grained attacks like GreedyPixel effectively manipulate diffusion model latent codes, whereas coarser black-box attacks (e.g., Square [44] and BruSLe [45]) fail. Comparing columns 6–10, increasing GreedyPixel’s query budget and sparsity allowance enhances model inversion accuracy by modifying more pixels but at a significantly higher computational cost.
V-C Out-Of-Domain Detection
We evaluate the proposed OODD framework on both face and non-face datasets across four face recognition models. As shown in Fig. 11, the model inversion outputs for out-of-domain inputs primarily fall into the two failure cases described in Sec. III, each of which serves as an out-of-domain indicator within our framework. The results in Tab. V demonstrate a low detection rate for face datasets (LFW and CelebA-HQ) and a high detection rate for non-face inputs (ImageNet), validating the effectiveness of our detection.

### v2

@@ 2.4.2 Face Detection via MTCNN
2.5 Step (b) - Top N Latent Code Selection
In Step (a), DiffUMI generates a set of $V$ randomly sampled yet reliable latent codes and their corresponding reconstructions $\{(x_{G_{v}},\hat{x}_{v})\}_{v=1}^{V}$. Instead of processing all $V$ candidates, it may select the one with the highest initial embedding similarity to the target, expecting lowest computational costs and distortions. However, the latent code yielding the best initial reconstruction may not always produce the optimal result after adversarial manipulation (Fig. 6). To address this, DiffUMI refines by choosing the top $N$ latent codes from $V$ (Fig. 16 of Appendix C), where increasing $N$ improves model inversion performance but incurs higher computational overhead.
@@ 2.5 Step (b) - Top N Latent Code Selection
In particular, the $V$ reliable pairs are input into the target model, which outputs $V$ feature embeddings. Note that $\hat{x}_{v}=G(x_{G_{v}})$, previously computed and stored during Step (a) for face detection (Sec. 2.4.2), is used. Similarities, as described in Eq. 6, are calculated between these $V$ embeddings and the target embedding. The top $N$ embeddings, exhibiting the highest similarity, are retained for Step (c) as the initial latent codes for adversarial manipulation.
In the black-box setting, as outlined in Sec. 2.2, each of the $(x_{G_{v}},\hat{x}_{v})$ pairs requires access to the target model, leading to a total of $V$ queries in Step (b):
@@
2.6 Step (c) - Latent Code Manipulation
@@
As outlined in Sec. 2.3, a randomly sampled latent code rarely meets the attack objective of reconstructing a facial image that matches the target identity, even with the top-1 selection from Step (b). Therefore, refining the initial latent code is crucial to align the reconstructed face with the target. Existing methods manipulate latent codes starting from either the top-1 code (its limitation was discussed in Sec. 2.5) or a fusion of the top $N$ selections [9]. However, we observe that fusion significantly reduces similarity compared to top-1 or top-2 selections. Moreover, relying on a single fused initialization may fail to ensure effective reconstruction, similar to the limitations of the top-1 approach. To address these issues, we propose the Ranked Adversary method, which retains the top $N$ selection strategy while prioritizing latent codes with higher similarity for earlier adversarial manipulation. This approach includes early termination once the attack objective is met, optimizing computational efficiency without sacrificing effectiveness.
2.6.1 Algorithm of Ranked Adversary
@@
The Ranked Adversary approach (Fig. 17 of Appendix C) begins the refinement process with the top-1 latent code $x_{G_{1}}$, applying an adversarial attack strategy to iteratively adjust $x_{G_{1}}$ in pursuit of the objective defined in Eq. 7, guided by the objective function in Eq. 6. The process concludes when the similarity measure $\mathcal{L}_{1}$ exceeds the predefined attack threshold $\tau_{A}$, indicating successful manipulation where the reconstructed image sufficiently matches the target. In such cases, the reconstructed image $\hat{x}_{1}=G(x_{G_{1}}+\delta_{1})$ is produced as the final model inversion result. If $\mathcal{L}_{1}$ remains below $\tau_{A}$ after the maximum adversarial iterations $t_{max}$, the top-2 latent code $x_{G_{2}}$ undergoes the same optimization process. If none of the top $N$ latent codes achieve $\mathcal{L}_{n}\geq\tau_{A},\ n=1,\dots,N$, the $\hat{x}_{n}$ with the highest $\mathcal{L}_{n}$ is selected as the final output.
@@
As shown in Fig. 7, setting $\tau_{A}=\tau_{F}$ may ensure successful target matching, but only on the target model. Increasing $\tau_{A}$ beyond $\tau_{F}$ enhances robustness, particularly on the test model. However, excessively high $\tau_{A}$ or the absence of early stopping may lead to the attack proceeds beyond the point of achieving $\mathcal{L}>0.98$, yielding negligible gains in optimization while exacerbating overfitting and diminishing generalization to the test model, as achieving such a threshold often forces adversarial manipulation to introduce artifacts that overfit the target model and exhaust all attack iterations. Therefore, we define $\tau_{A}$ for attacking the specific model as the maximum embedding similarity achievable by real facial images in this model, which represents the best case where without any overfitting artifacts, the model can achieve the best similarity within the same identity. To compute this, we feed the entire face dataset $\mathcal{X}_{real}$ (all real faces) into the target model, then compute the maximum similarity between any two images. The similarity between different identities is usually lower than that between the same identity. Thus, we define $\tau_{A}$ as follows:
@@
2.6.2 Adversarial Attack
Successful manipulation of unconditional diffusion generation does not require a new adversarial attack backbone, but instead relies on a fine-grained algorithm [38]. Ranked Adversary employs APGD [36] in the white-box setting and GreedyPixel [38] in the black-box setting. While various adversarial attack algorithms can be adapted for latent code manipulation within our framework, our empirical observation shows that only these fine-grained methods yield satisfactory reconstructions without introducing significant distortions or artifacts. This is due to the sensitivity of the diffusion model’s latent space, where even small perturbations can notably degrade reconstruction quality.
Additionally, we evaluated two alternative black-box attacks: Square attack [47], which uses the same $L_{2}$-norm constrained perturbation magnitude $\epsilon$ as APGD, and BruSLe attack [48], a pixel-wise attack like GreedyPixel that enforces sparsity constraints rather than directly constraining $\epsilon$. However, both methods performed inferiorly compared to GreedyPixel, as detailed in Sec. 5.2.
2.6.3 Query Efficiency in the Black-Box Setting
@@
3 Out-Of-Domain Detection
Deep learning models map both in-domain and out-of-domain inputs to the same feature space, making it challenging to distinguish between them based solely on embeddings. For example, in a face recognition system, both a human face and a non-face object (e.g., a cat) produce embeddings with identical dimensions and numerical ranges, despite their semantic differences. To address this, we leverage the high-fidelity reconstruction capabilities of our model inversion framework to develop an application of OODD that identifies out-of-domain inputs based on their embeddings. As illustrated in Fig. 8, two common failure cases arise in model inversion, sometimes concurrently. In-domain inputs, such as facial images, typically yield accurate, recognizable reconstructions. In contrast, out-of-domain inputs are more likely to exhibit at least one of these failure cases. Hence, we classify an input as out-of-domain if either failure is detected.
@@ 3 Out-Of-Domain Detection
where $\tau_{F_{m}}$ is the predefined similarity threshold for the minimum equal error rate in face recognition.
@@
  • 1 A higher detection confidence threshold is applied during latent code generation to ensure reliable initialization, as generating a random Gaussian template is computationally efficient. All parameters refer to Tab. XIV in Appendix C for notation definitions.
@@
4 Experimental Settings
Face Recognition Models. We assess model inversion attacks on two widely used models, FaceNet [4] and ArcFace [5], and two privacy-preserving models, DCTDP [6] and PartialFace [7], which are designed to mitigate inversion attacks. The decision thresholds for all models are set at the minimum equal error rate, as outlined in Tab. II.
@@ 4 Experimental Settings
Benchmark. We benchmark DiffUMI against MAP2V [9] in the white-box setting. As shown in Tab. I, MAP2V is the most recent and most relevant state-of-the-art baseline, sharing our training-free setup and focus on open-set face recognition. While DiffUMI also outperforms training-dependent attacks [27, 17], as detailed in Appendix A, they are not our primary baseline due to their limitation that each target model requires a separately trained generator, resulting in substantial computational overhead.
Experimental Setup. The configurations for the white-box DiffUMI attack and OODD are provided in Tab. II, while the settings for the black-box DiffUMI attack are outlined in Fig. 11. Specifically, $\tau_{F}$ is the cosine similarity threshold in face recognition, optimized to minimize the equal error rate. $\tau_{K}$ and $\tau_{D}$ in latent code generation are set close to the upper bound (1.0) to ensure highly reliable initialization. $\tau_{A}$ is defined in Eq. 14. For OODD, $\tau_{D}$ is set to maintain a maximum false positive rate of 5% for in-domain inputs. In the ablation studies, we vary one parameter at a time while holding all others constant. Experiments are conducted on an NVIDIA A100 40GB GPU.
Evaluation Protocols. Model inversion attacks and OODD are executed directly on the target model, often leading to outputs that overfit this model. Consequently, exceptional performance on the target model may stem from adversarial artifacts rather than the accurate reconstruction of facial features, a scenario classified as a failure in Sec. 3. To thoroughly evaluate the effectiveness of privacy attacks and OODD, we adopt the following evaluation protocols.
Figure 9: A challenging example of confusing choices in questionnaire design for identifying target identities from reconstructions. We provided two photos for each of three target options, with the order randomized. Some participants remarked that the two male choices (the first and second targets) looked too similar, making them appear to be the same person.
User Study: Following the attack objectives defined in Sec. 2.2, we asked ten participants to evaluate three groups of images, specifically, to identify the target identities based on reconstructions and to judge whether each reconstruction depicts the same person as the target identity. Each evaluation included 60 images in total, drawn from two datasets. To make the questionnaire sufficiently challenging, we introduced confusing choices by selecting identities whose facial embeddings have a similarity of approximately $80\%\times\tau_{F}$ to the target identity (as shown in Fig. 9), where a similarity of $\geq\tau_{F}$ is considered a match.
@@
$\text{Type II}=\quad\frac{\sum_{i=1}^{I}\sum_{j=1}^{J}\mathbb{I}\left(S(F(\hat{x_{i}}),F(x^{j\neq T}_{i}))\geq\tau_{F}\right)}{I},$ (22)
@@
For further validation, we also report Type I accuracy results in Appendix D, which reflect the strength of the inversion attack and highlight potential privacy vulnerabilities.
OODD Rate: The OODD rate quantifies the failure rate of model inversion, based on the two failure cases outlined in Sec. 3. For in-domain inputs, a lower OODD rate is preferable, indicating fewer inversion failures. Conversely, for out-of-domain inputs, a higher OODD rate is desirable, reflecting improved detection performance.
Joint Evaluation Using Target and Test Models: Test models, as defined in Sec. 3, refer to those that differ from the target model. Evaluating on these models helps mitigate overfitting to the target model. In this study, four face recognition models are utilized, with one designated as the target model and the remaining three as test models for each attack. However, testing exclusively on test models may introduce bias, as the test models vary for each target model, leading to an unfair comparison of model vulnerability to privacy risks. To address this, we adopt a joint evaluation approach that includes both the target and test models, ensuring consistent evaluation across all models. This approach mitigates the impact of overfitting to the target model alone.
Note that OODD is evaluated exclusively on test models, as detailed in Sec. 3. This is because out-of-domain inputs can only be matched with their reconstruction (similar to Type I accuracy), and such matching is always successful on the target model due to overfitting.
5 Performance
5.1 White-Box Model Inverison
We assess the vulnerability of face recognition models to privacy threats using DiffUMI in the white-box setting. As shown in Tab. III, all four models, including privacy-preserving variants, fail to prevent privacy leakage. DiffUMI achieves Type II accuracy between 74.51% and 94.72% across all models and datasets. Notably, the oldest standard model, FaceNet, demonstrates the highest resistance, highlighting the limitations of existing privacy-preserving techniques. Furthermore, the user study results in Tab. IV reflect a high success rate of our privacy attack, as confirmed through direct human evaluation.
@@ 5.1 White-Box Model Inverison
TABLE IV: User study results. Target Model is the privacy-preserving model, Partialface [7].
      Question       Dataset       Accuracy (%) $\uparrow$
      Find target       CelebA       80.0 (16.0 / 20.0)
      Find target       LFW       79.5 (15.9 / 20.0)
      Same person?       LFW       73.0 (14.6 / 20.0)
Compared to the benchmark, as shown in Tab. V, DiffUMI consistently surpasses MAP2V in Type II accuracy and achieves similarity values closer to the attack threshold $\tau_{A}$ across all scenarios. The “random” rows represent unguided diffusion model generations, which fail to match targets, emphasizing the necessity of a structured inversion approach. The visualization in Fig. 10 further highlights DiffUMI’s superiority, demonstrating higher identity recovery accuracy than MAP2V. Our approach produces full headshot-style reconstructions, achieving optimal granularity for identity recovery.
TABLE V: Performance of our DiffUMI across four face recognition models, compared to the benchmark attack MAP2V [9].
@@
  • 3 Average Type II accuracy across four test models [5, 4, 6, 7], as per the joint evaluation using target and test models outlined in Sec. 4.
@@
Figure 11: Performance of DiffUMI in the Black-Box Setting. Black-box DiffUMI achieves slight lower reconstruction accuracy to its white-box counterpart (columns 3 vs. 10) in terms of Type II accuracy. Among black-box attack methods (columns 4-6), the fine-grained GreedyPixel algorithm, which introduces adversarial patterns akin to white-box attacks, is the only viable approach for manipulating diffusion model latent codes. Comparing columns 6–10, increasing the attack budget (query or sparsity) enhances model inversion accuracy. The target model is the privacy-preserving PartialFace [7], and the test model is ArcFace [5].
5.2 Black-Box Model Inversion
We evaluate DiffUMI in the black-box setting using various adversarial attack algorithms and compare its performance to the white-box counterpart. As shown in Fig. 11 (columns 3 vs. 10), black-box DiffUMI (GreedyPixel) achieves Type II accuracy slightly lower than its white-box version (APGD) but incurs substantially higher computational costs in queries. Moreover, Fig. 11 (columns 4–6) demonstrates that only fine-grained attacks like GreedyPixel effectively manipulate diffusion model latent codes, whereas coarser black-box attacks (e.g., Square [47] and BruSLe [48]) fail. Comparing columns 6–10, increasing GreedyPixel’s query budget and sparsity allowance enhances model inversion accuracy by modifying more pixels.
5.3 Out-Of-Domain Detection
We evaluate the proposed OODD framework on both face and non-face datasets across four models. As shown in Fig. 12, the model inversion outputs for out-of-domain inputs primarily fall into the two failure cases described in Sec. 3, each of which serves as an out-of-domain indicator within our framework. The results in Tab. VI demonstrate a low detection rate for face datasets (LFW and CelebA-HQ) and a high detection rate for non-face inputs (ImageNet), validating the effectiveness of our detection.